# Generalized discrete Markov spectra

### v2

@@ 7. Formulation of Generalized Uniqueness Conjecture
This conjecture is equivalent to the following (see [aig]*Corollary 9.30):
@@
As a generalization of Conjecture 7.1, the author and Matsushita [gyomatsu] posed the following question:
@@
A counterexample of Question 7.3 arises when $k_{1},k_{2},k_{3}$ are distinct ([gyomatsu]*Remark 10). In fact, both $(1,81,17)$ and $(7,81,2)$ are $(1,2,0)$-GM triples (and moreover positive integer solutions of the $(1,2,0)$-GM equation).
According to [hurwitz], Question 7.4 has the negative answer for $\mathcal{M}\setminus\mathcal{M}_{0,0,0}$. Here we shall introduce a concrete example of this phenomenon. One sees that when $k_{1},k_{2},k_{3}$ are not all zero there is an obvious counterexample coming from Theorem 6.1. For a generalized strongly admissible sequence $s(t)$, set $\alpha=[s(t)^{\infty}]$ and $\beta=[s^{*}(\frac{1}{t})^{\infty}]$. By Theorems 2.20 and 6.1, we have $\mathcal{L}(\alpha)=\mathcal{L}(\beta)$. By Remark 4.5 (7), the periodic parts of $s(t)^{\infty}$ and $s^{*}(\frac{1}{t})^{\infty}$ are reverse-ordered with respect to each other, and in general they do not coincide unless $k_{1}=k_{2}=k_{3}=0$. By Serret’s theorem, two irrationals are $GL(2,\mathbb{Z})$-equivalent if and only if their periodic parts coincide, so $\alpha$ and $\beta$ need not be $GL(2,\mathbb{Z})$-equivalent. Moreover, there exist counterexamples that do not fall into this case. Such counterexample can be constructed using the counterexample of Question 7.3. To see this, set
@@
Conjecture 7.5 ([aig]*Uniqueness conjecture IV).
@@
Theorem 7.7 ([mcshane, llrs, banaian, banaian-sen]).

### v3

@@ 7. Formulation of Generalized Uniqueness Conjecture
This conjecture is equivalent to the following (see [1]*Corollary 9.30):
@@
As a generalization of Conjecture 7.1, the author and Matsushita [18] posed the following question:
@@
A counterexample of Question 7.3 arises when $k_{1},k_{2},k_{3}$ are distinct ([18]*Remark 10). In fact, both $(1,81,17)$ and $(7,81,2)$ are $(1,2,0)$-GM triples (and moreover positive integer solutions of the $(1,2,0)$-GM equation).
According to [20], Question 7.4 has the negative answer for $\mathcal{M}\setminus\mathcal{M}_{0,0,0}$. Here we shall introduce a concrete example of this phenomenon. One sees that when $k_{1},k_{2},k_{3}$ are not all zero there is an obvious counterexample coming from Theorem 6.1. For a generalized strongly admissible sequence $s(t)$, set $\alpha=[s(t)^{\infty}]$ and $\beta=[s^{*}(\frac{1}{t})^{\infty}]$. By Theorems 2.20 and 6.1, we have $\mathcal{L}(\alpha)=\mathcal{L}(\beta)$. By Remark 4.5 (7), the periodic parts of $s(t)^{\infty}$ and $s^{*}(\frac{1}{t})^{\infty}$ are reverse-ordered with respect to each other, and in general they do not coincide unless $k_{1}=k_{2}=k_{3}=0$. By Serret’s theorem, two irrationals are $GL(2,\mathbb{Z})$-equivalent if and only if their periodic parts coincide, so $\alpha$ and $\beta$ need not be $GL(2,\mathbb{Z})$-equivalent. Moreover, there exist counterexamples that do not fall into this case. Such counterexample can be constructed using the counterexample of Question 7.3. To see this, set
@@
Conjecture 7.5 ([1]*Uniqueness conjecture IV).
@@
Theorem 7.7 ([25, 21, 5, 3]).